Ritonavir is an anti-viral drug given to patients with HIV. While the normal recommended adult dose is 600 mg twice daily, the recommended dose for children over 1 month old is 350 mg/m2. What dose should be given to a 9 month old infant that is 27 inches in height and 18 lbs? Ritonavir is available 80 mg/mL. How many mL should be given to the infant?

Label all numbers with its corresponding unit and drug

Place zeros in front of decimals (0.1)
Answer #1

ANSWER.

* Calculation of volume(ml) and dose(mg) of drug required for this infant.

              Recommended dose for children = 350 mg/m2 (over 1 month old).

        Body surface area (m2) of a 9 month old infant.

               Formula

                  Body surface area = square root ((height * weight) / 3600)

                       Weight of the infant = 18 lbs.

                Convert lbs to kg for calculation

                                1 lbs = 0.45 kg

                So

                     Weight of the infant = 18 * 0.45

                                                     = 8.1 kg.

                      Height of the infant = 27 inch

               Convert inch to cm for calculation

                                 1 inch = 2.5 cm

             So

                    Height of the infant = 27 * 2.5

                                                   = 67.5 cm.

      As per above formula for BSA

          Body surface area of the infant = square root ( 8.1 * 67.5) / 3600)

                                                 = square root (546.8 / 3600)

                                                 = square root ( 0.15)

                                                 = 0.38 m2.

     * Dose required for this 9 month old infant

                                           350 mg = m2

                                         X mg = 0.38 m2

                                                X mg = 350 * 0.38

                                                         = 133 mg of ritonavir

   * Volume(ml) of drug required for this infant.

                 Formula = Desire dose/Dose in hand * Quantity.

          Here

                 Desired dose(D) = 133 mg.

                 Dose in hand (H) = 80 mg.

                     Quantity (Q) = 1 ml.

      As per above formula

             Volume(ml) of Ritonavir required = D/H * Q

                                                           = 133 / 80 * 1

                                                           = 1.7 ml of ritonavir.